jayce bought 2 bath towels and returned 3 hand towels. his sister jayna bought 3 bath towels and 3 hand towels. jayces bill was $5. jaynas bill was $45. what are the prices of a bath towel and a hand towel
Answer by josgarithmetic(39616)

b, price of Bath towels
h, price of Hand towels
-
Jayce: 2b-3h=5
Jayna: 3b+3h=45
-
THAT is the system to solve. Notice subtraction of 3h in Jayce's bill because of RETURNING some towels for credit.
Add the two equations for immediate use of Elimination Method to first solve for b.
5b=50
b=10----One answer
-
2b-5=3h
h=(2b-5)/3
h=(2*10-5)/3
h=15/3
h=5-----Other answer